OpenStudy (anonymous):

find the exact value of tan(pi/12)

OpenStudy (anonymous):

\[\tan(x-y) = \frac{tanx-tany}{1+tanxtany}\]

OpenStudy (anonymous):

\[\tan(\dfrac{\pi}{12}) = \tan( \frac{4\pi}{12}-\frac{3\pi}{12}) \]

OpenStudy (anonymous):

\[\tan(\frac{\pi}{3}-\frac{\pi}{4})\]

OpenStudy (anonymous):

\[\tan(\dfrac{\pi}{12}) =\frac{ \tan(\frac{ \pi }{ 3 }) - \tan(\frac{ \pi }{ 4 }) }{ 1-\tan(\frac{ \pi }{ 3 })(\tan(\frac{ \pi }{ 4 }) } \]